# Global graph features unveiled by unsupervised geometric deep learning

### v1

@@ Abstract
Graphs provide a powerful framework for modeling complex systems, but their structural variability makes analysis and classification challenging. To address this, we introduce GAUDI (Graph Autoencoder Uncovering Descriptive Information), a novel unsupervised geometric deep learning framework that captures both local details and global structure. GAUDI employs an innovative hourglass architecture with hierarchical pooling and upsampling layers, linked through skip connections to preserve essential connectivity information throughout the encoding–decoding process. By mapping different realizations of a system — generated from the same underlying parameters — into a continuous, structured latent space, GAUDI disentangles invariant process-level features from stochastic noise. We demonstrate its power across multiple applications, including modeling small-world networks, characterizing protein assemblies from super-resolution microscopy, analyzing collective motion in the Vicsek model, and capturing age-related changes in brain connectivity. This approach not only improves the analysis of complex graphs but also provides new insights into emergent phenomena across diverse scientific domains.
@@ Introduction
More recently, the performance of GNNs in graph unsupervised learning has been advanced through graph contrastive learning [13]. Graph contrastive learning uses augmentations to create multiple versions of the graph by introducing slight variations but retaining the essential structure and properties. It thus maximizes the similarity between different views of the same graph and minimizes the similarity between views of different graphs. In this way, graph contrastive learning obtains generalizable graph representations for various downstream tasks [14]. A major limitation of most existing graph contrastive learning methods is that they derive node representations primarily from local neighborhood information and have inherent limitations in capturing global knowledge, either structural or semantic [15]. In addition, their predominant focus on instance-level contrast leads to a latent space where inputs with different augmentations have similar representations, but the underlying semantic structure of the input graph is largely ignored [16].
@@
In this article, we propose a Graph Autoencoder Uncovering Descriptive Information (GAUDI), a geometric-deep learning framework that captures multiple essential features and properties of complex systems represented as graphs. By combining graph convolution with MinCut pooling and upsampling, GAUDI gradually reduces the size of a graph to a compressed version in a low-dimensional latent space. Connectivity information is directly passed from the encoder to the decoder so that the latent space can capture the relevant parameters and global structural characteristics of complex systems. In the latent space, realizations of a system with the same underlying parameters are placed close together, assembling a structure that corresponds to the inherent parameter space. The latent space representation is subsequently decoded to reconstruct the input graph, enabling unsupervised training of the model. We demonstrate GAUDI’s capability to uncover descriptive parameters of graph-represented systems across four diverse scenarios: Watts-Strogatz small-world graphs [5], protein assemblies from single-molecule localization microscopy [17], collective behavior in active matter [18], and brain connectivity [19]. By uncovering the parameters describing these systems, GAUDI opens a path for a deeper understanding of their underlying complexity and aids in identifying the similarities and disparities among different conditions.
@@ Results
The objective of GAUDI is to map different realizations of the state of a complex system – described by the same values of the underlying parameters – into nearby points on the latent space in an unsupervised manner, as schematically shown in Fig. 1a. The realization of the system is represented through a graph, where nodes correspond to its elements and edges denote interactions and relationships. Each node is characterized by a set of features that reflect the state of the element within the complex system. Graph connectivity is encoded by the adjacency matrix, which indicates which pairs of nodes are connected.
@@
Through self-supervised training based on graph reconstruction, GAUDI learns to align the node features of the reconstructed graph with those of the input graph, thus forcing the latent space to be highly expressive.
@@
Figure 2: GAUDI representation of Watts-Strogatz small-world graphs in latent space. GAUDI encodes Watts-Strogatz graphs, characterized by node degree $C$ and rewiring probability $p$, into latent variables. a, These graphs model structures ranging from highly ordered to completely random as $p$ increases. As $C$ increases, nodes form more connections, resulting in higher overall connectivity. b, Scatter plot of the first two principal components of the graphs’ latent representations, with each point representing a compressed graph. The shape of the scatter points indicates the node degree $C$, where GAUDI effectively clusters graphs based on this parameter. In contrast, the color gradient represents the rewiring probability $p$, showing a smooth transition without distinct clustering, thus illustrating the seamless integration of $p$ variations within the latent space. The inset shows the latent space for the Watts-Strogatz graphs obtained using graph contrastive learning.
@@
To compare GAUDI with unsupervised graph contrastive learning, we applied the method proposed by Wang et al. [28] to the same Watts-Strogatz small-world graphs with slight modifications to the original architecture and hyperparameters. Specifically, we replaced the embedding layer with a linear layer to account for the differences in node and edge features between the Watts-Strogatz graphs and the molecular graphs used in the original work. Additionally, we reduced the latent space dimension to 8 to match the one used for GAUDI. During training, input graphs were duplicated, nodes were randomly masked, and a contrastive loss was utilized to bring graphs originating from the same input graph closer together in the latent space.
As shown in the inset of Fig. 2b, the latent space obtained using the graph contrastive learning captures local similarities between the graphs by placing similar graphs close to each other. However, there is no global structure encoded in the latent space, making it impossible to determine the parameters of a graph from its latent space representation. Although similar graphs are grouped together, very different graphs can also be placed close to each other.
In contrast, as seen in the main plot of Fig. 2b, GAUDI effectively captures the parameters of the graphs both locally and globally. Locally, it places similar graphs close together, ensuring that graphs with similar characteristics are grouped. Globally, it maintains a clear structure in the latent space, making it easy to distinguish graphs based on their parameters. This dual capability marks a significant improvement over graph contrastive learning, which fails to encode a coherent global structure in the latent space.
@@ Methods
[…]
@@
with .$\left\lVert.\right\rVert_{F}$ denoting the Frobenius norm, $D$ denoting the degree matrix of $A$, and $\nu$ denoting the number of nodes of the pooled graph.
@@
Watts-Strogatz small-world graphs are composed by nodes that placed in a ring and being connected in a manner provided by the two parameters $C$, the number of neighbouring nodes, and $p$, the rewiring probability [5]. Here, we use 80 nodes for each graph. To begin with, each node is connected to it’s $C$ nearest neighbors, where $C$ denotes an even number. Here, we use the values 2, 4, 6, and 8 for $C$. From this representation each edge is rewired once with the probability $p$, ranging from 00 to $1$. This rewiring removes the existing edge, and connects the node to a randomly assigned node. This results in a regular graph for $p$ equal or close to zero, while an increasing $p$ leads to a more random graph (Fig. 2a).
@@
To form a graph from a sample, the data points are seen as nodes, and each node is connected to its six most distant neighbors within that sample. The graphs are normalized by centering the coordinates of the nodes around (0,0) and by dividing the coordinates by the standard deviation of all coordinates of the graph. The node features of these graphs are set to be the inverse distance of their 2-dimensional coordinate to the center of the graph, while the edge feature used is the inverse length of the edge.
@@
In each simulation, we use $N=100$ particles within a square with a side length equal to 100 units. The particles maintain a constant speed set to $v=1$ unit per time step, and the time step is defined as $\Delta t=1$. The orientation of particle $i$ at time step $t+1$ is updated using the following equation: $\theta_{i,t+1}=\langle\theta_{j,t}\rangle+\eta\cdot W_{i,t}\cdot\Delta t$, where particles $j$ have a distance smaller than the flocking radius $R_{\rm f}$ to particle $i$ ($|\mathbf{r}_{j,t}-\mathbf{r}_{i,t}|<R_{\rm f}$), and $W_{i,t}$ are random numbers from the uniform distribution $[-\frac{1}{2},\frac{1}{2}]$. Once the orientations have been updated, the positions are adjusted: $\mathbf{r}_{i,t+1}=\mathbf{r}_{i,t}+\mathbf{v}\cdot\Delta t$, with the velocity $\mathbf{v}=v\cdot(cos(\theta_{i,t})\hat{x}+sin(\theta_{i,t})\hat{y})$

### v2

@@ Abstract
Graphs provide a powerful framework for modeling complex systems, but their structural variability poses significant challenges for analysis and classification. To address these challenges, we introduce GAUDI (Graph Autoencoder Uncovering Descriptive Information), a novel unsupervised geometric deep learning framework designed to capture both local details and global structure. GAUDI employs an innovative hourglass architecture with hierarchical pooling and upsampling layers linked through skip connections, which preserve essential connectivity information throughout the encoding–decoding process. Even though identical or highly similar underlying parameters describing a system’s state can lead to significant variability in graph realizations, GAUDI consistently maps them into nearby regions of a structured and continuous latent space, effectively disentangling invariant process-level features from stochastic noise. We demonstrate GAUDI’s versatility across multiple applications, including small-world networks modeling, characterization of protein assemblies from super-resolution microscopy, analysis of collective motion in the Vicsek model, and identification of age-related changes in brain connectivity. Comparison with related approaches highlights GAUDI’s superior performance in analyzing complex graphs, providing new insights into emergent phenomena across diverse scientific domains.
@@ Introduction
In parallel, the performance of GNNs in graph unsupervised learning has been advanced through graph contrastive learning [13]. Graph contrastive learning uses augmentations to create multiple versions of the graph by introducing slight variations but retaining the essential structure and properties. It thus maximizes the similarity between augmented views of the same graph and minimizes the similarity between representations of different graphs. In this way, graph contrastive learning obtains generalizable graph representations for various downstream tasks [14].
A major limitation of most existing graph contrastive learning methods is that they derive node representations primarily from local neighborhood information and have inherent limitations in capturing global knowledge, either structural or semantic [15]. In addition, their predominant focus on instance-level contrast leads to a latent space where inputs with different augmentations have similar representations, but the underlying semantic structure of the input graph is largely ignored [16].
@@
In this article, we propose a Graph Autoencoder Uncovering Descriptive Information (GAUDI), a geometric-deep learning framework that captures multiple essential features and properties of complex systems represented as graphs. By combining graph convolution with MinCut pooling and upsampling, GAUDI gradually reduces the size of a graph to a compressed version in a low-dimensional latent space. Connectivity information is directly passed from the encoder to the decoder so that the latent space can capture the relevant parameters and global structural characteristics of complex systems. In the latent space, realizations of a system with the same underlying parameters are placed close together, assembling a structure that corresponds to the inherent parameter space. The latent space representation is subsequently decoded to reconstruct the input graph, enabling unsupervised training of the model.
We demonstrate GAUDI’s capability to uncover descriptive parameters of graph-represented systems across four diverse scenarios: Watts-Strogatz small-world graphs [5], protein assemblies from single-molecule localization microscopy [17], collective behavior in active matter [18], and brain connectivity [19]. In each case, GAUDI effectively maps system realizations into a structured latent space that reflects underlying physical or biological parameters. By doing so, it facilitates a deeper understanding of the system’s complexity and enables meaningful comparisons between conditions. We benchmarked against established unsupervised graph learning methods, including autoencoders [20] and contrastive learning approaches [21], GAUDI achieve superior performance in disentangling relevant features and capturing both local and global graph properties.
@@ Results
The objective of GAUDI is to map different realizations of the state of a complex system — described by the same values of the underlying parameters — into nearby points on the latent space in an unsupervised manner, as schematically shown in Fig. 1a. The realization of the system is represented through a graph, where nodes correspond to its elements and edges denote interactions and relationships. Each node is characterized by a set of features that reflect the state of the element within the complex system. Graph connectivity is encoded by the adjacency matrix, which indicates which pairs of nodes are connected.
@@
Through self-supervised training based on graph-to-graph reconstruction, GAUDI learns to align the node features of the reconstructed graph with those of the input graph, thus forcing the latent space to be highly expressive.
@@
Figure 2: GAUDI representation of Watts-Strogatz small-world graphs in latent space. GAUDI encodes Watts-Strogatz graphs, characterized by node degree $C$ and rewiring probability $p$, into latent variables. a, These graphs model structures ranging from highly ordered to completely random as $p$ increases. As $C$ increases, nodes form more connections, resulting in higher overall connectivity. b, Scatter plot of the first two principal components of the graphs’ latent representations, with each point representing a compressed graph. The shape of the scatter points indicates the node degree $C$, where GAUDI effectively clusters graphs based on this parameter. In contrast, the color gradient represents the rewiring probability $p$, showing a smooth transition without distinct clustering, thus illustrating the seamless integration of $p$ variations within the latent space.
@@
GAUDI outperforms existing methods across diverse tasks
To benchmark GAUDI against state-of-the-art methods for unsupervised graph representation learning, we compare its performance with two alternative approaches: a graph contrastive learning approach [21] and a graph autoencoder [20].
We selected these benchmarks as they represent prominent and conceptually distinct approaches in unsupervised graph learning. Graph contrastive learning with structural augmentations is a widely used strategy for learning graph-level representations, while autoencoders have recently been designed to capture both local node-level and global graph features. Moreover, these methods generate a single latent representation per graph, making them directly comparable to GAUDI without requiring substantial architectural modifications. We evaluate all models consistently across the four graph-represented systems studied in this work.
We begin with a qualitative analysis by visualizing the first two principal components of the latent representations, as shown in Fig. 6. Across all examples, GAUDI effectively captures both local and global aspects of the graphs’ underlying parameters (Fig. 6a). Locally, it places similar graphs near each other, forming tight clusters based on graph characteristics. Globally, it organizes the latent space in a structured and interpretable way, such that the graph-generating parameters are reflected in the global layout of the embeddings.
[…]
Figure 6: Comparison of latent space representations obtained with different methods. The first two principal components of the latent representations are shown for each method (rows) and dataset (columns): (a) GAUDI reproduced from Figs. 2b, 3c, 4b, and 5c; (b) graph contrastive learning with node and edge masking (CL-n); (c) contrastive learning with subgraph removal (CL-s); (d) MIAGAE with 8 latent dimensions (MIAGAE8); and (e) MIAGAE with 64 latent dimensions (MIAGAE64). Each column corresponds to one of the datasets analyzed in this study: Watts–Strogatz, SMLM, Vicsek, and brain connectivity graphs. Each point represents a graph, with marker shape and color indicating its generating parameters.
For contrastive learning, we implemented the method by Wang et al. [21], which augments graphs by applying node and edge masking or subgraph removal, and optimizes a contrastive loss that brings augmentations of the same input graph closer in latent space. To adapt the method to our data, we introduced minor architectural and hyperparameter modifications. Specifically, we replaced the input embedding layer with a linear layer to accommodate our node and edge features, and set the latent dimension to 8 to ensure fair comparison. We evaluated two variants: node and edge masking (CL-n, 25% masked, Fig. 6b) and subgraph removal (CL-s, 25% removed nodes and edges, Fig. 6c). Both CL-n and CL-s group similar graphs locally, but fail to maintain a meaningful global organization in the latent space. This results in clusters of similar graphs, but with dissimilar graphs also frequently embedded nearby, making it difficult to infer parameters from the representation.
To compare GAUDI with a graph autoencoder, we used the Multi-kernel Inductive Attention Graph Autoencoder (MIAGAE) by Ge et al. [20]. MIAGAE reduces graph size via multi-kernel inductive convolutions and attention-based pooling in the encoder, and reconstructs it through inductive upsampling layers. It is trained using the mean squared error between the original and reconstructed node features. Since MIAGAE does not aggregate graphs to a single node, we average the encoded node features to obtain a single latent vector per graph. We tested both a reduced version (MIAGAE-8, Fig. 6d) constrained to a 8-dimensional latent space for comparability, and the default version (MIAGAE-64, Fig. 6e) with 64 dimensions. Both MIAGAE variants preserve some local structure but lack a consistent global organization. While some similar graphs are placed together, marker shapes indicating categorical variables are intermixed, and color gradients reflecting continuous parameters lack a coherent trend. This limits the interpretability of the embeddings and impairs the extraction of generative parameters.
Dataset Method 1-NN accuracy Isomap gradient continuity AUC $R^{2}$
Watts-Strogatz GAUDI 1.00 0.99 - -
CL-n 0.51 0.24 - -
CL-s 0.61 0.22 - -
MIAGAE8 0.37 0.28 - -
MIAGAE64 0.50 0.77 - -
SMLM GAUDI 0.80 - 0.92 (0.94*) -
CL-n 0.78 - 0.84 (0.86*) -
CL-s 0.61 - 0.61 (0.76*) -
MIAGAE8 0.65 - 0.83 (0.86*) -
MIAGAE64 0.75 - 0.90 -
Vicsek GAUDI 0.93 0.98 - -
CL-n 0.49 0.88 - -
CL-s 0.75 0.95 - -
MIAGAE8 0.80 0.82 - -
MIAGAE64 0.79 0.82 - -
Brain connectivity GAUDI - 0.70 0.87 (0.87*) 0.51 (0.54*)
CL-n - 0.14 0.56 (0.63*) 0.04 (0.12*)
CL-s - 0.16 0.58 (0.65*) 0.05 (0.08*)
MIAGAE8 - 0.08 0.74 (0.74*) 0.27 (0.31*)
MIAGAE64 - 0.52 0.78 0.36
Table 1: Quantitative comparison of GAUDI, graph contrastive learning, and MIAGAE. We report the 1-Nearest Neighbor (1-NN) accuracy, class-wise Isomap gradient continuity score, area under the ROC curve (AUC), and coefficient of determination ($R^{2}$) for each method on four datasets: Watts–Strogatz small-world graphs, single-molecule localization microscopy (SMLM), Vicsek model simulations, and brain connectivity graphs from the CamCAN dataset. Note that not all metrics are applicable to all datasets, as their computation depends on the presence of discrete labels or continuous generative parameters. Contrastive learning is evaluated with node and edge masking (CL-n) and subgraph removal (CL-s). MIAGAE is tested with 8 (MIAGAE8) and 64 (MIAGAE64) latent dimensions. All metrics are computed from the first two principal components (PC1 and PC2) of the latent space, except where noted (*), in which case all 8 latent dimensions are used. Higher scores indicate better performance. Bold values indicate the best performance for each metric within each dataset.
To perform a quantitative comparison of the methods’ performance, we computed several metrics, summarized in Table 1. First, we use the 1-Nearest Neighbor (1-NN) accuracy in the PCA plane to measure whether graphs from the same class form local clusters while remaining distinct from others. Next, to assess whether continuous parameters vary smoothly within classes, we compute a class-wise Isomap gradient continuity score [42], defined as the mean absolute Spearman rank-order correlation [43, 44] between the Isomap coordinates and the continuous parameter of interest. For binary classification tasks, we report the area under the ROC curve (AUC) from a support vector machine trained on the latent space. Finally, to quantify how well the latent representations explain variation in continuous variables, we compute the coefficient of determination ($R^{2}$). Note that some metrics are dataset-specific, as their applicability depends on the presence of either continuous generative parameters or discrete class labels. Full metric details are provided in Metrics for performance evaluation.
As shown in Table 1, GAUDI consistently outperforms both graph contrastive learning and MIAGAE across all metrics and datasets. In particular, it achieves near-perfect class-wise Isomap gradient continuity scores for the Watts–Strogatz (0.99) and Vicsek (0.98) graphs, indicating an almost monotonic embedding of the continuous parameters $p$ and $\eta$ within the respective classes. For the SMLM data, GAUDI achieves the highest AUC (0.92 using the first two principal components and 0.94 when using all 8 latent dimensions), outperforming all other methods in this binary classification task. GAUDI also yields the highest continuity score (0.70) on the brain connectivity data and the highest $R^{2}$ value (0.51), further demonstrating its ability to uncover structure in more complex and noisy biological graphs. This strong global organization, combined with accurate local clustering, highlights GAUDI’s ability to simultaneously capture discrete structure and continuous variation—something that neither contrastive learning nor MIAGAE achieve effectively across all datasets.
@@ Discussion
Compared to state-of-the-art methods such as graph contrastive learning and MIAGAE, GAUDI consistently achieves superior performance across a range of datasets and metrics. Unlike contrastive learning approaches and other autoencoders, which often fail to organize the latent space globally — resulting in poor interpretability of continuous parameters — GAUDI produces structured and disentangled latent spaces that reflect both categorical and continuous generative variables. GAUDI’s ability to preserve monotonic trends in continuous parameters and maintain separability across classes demonstrates its effectiveness in encoding both local and global structure, even in the presence of noise or intrinsic stochasticity.
@@ Methods
where $\mathbf{m}_{ij}$ denotes the message passed from node $j$ to node $i$, $\left[,\right]$ represents concatenation, $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ are the feature vectors of nodes $i$ and $j$, respectively, $\mathbf{e}_{ij}$ denotes the edge features of the edge between nodes $i$ and $j$, and the function $\gamma$ is a fully-connected dense layer followed by a ReLU activation. Subsequently, the features of each node are updated by aggregating the messages $\mathbf{m}_{ij}$ received from its neighbors, as follows,
@@
with $\left\lVert.\right\rVert_{F}$ denoting the Frobenius norm, $D$ denoting the degree matrix of $A$, and $\nu$ denoting the number of nodes of the pooled graph.
@@
Watts-Strogatz small-world graphs are composed by nodes that placed in a ring and being connected in a manner provided by the two parameters $C$, the number of neighboring nodes, and $p$, the rewiring probability [5]. Here, we use 80 nodes for each graph. To begin with, each node is connected to it’s $C$ nearest neighbors, where $C$ denotes an even number. Here, we use the values 2, 4, 6, and 8 for $C$. From this representation each edge is rewired once with the probability $p$, ranging from 0 to $1$. This rewiring removes the existing edge, and connects the node to a randomly assigned node. This results in a regular graph for $p$ equal or close to zero, while an increasing $p$ leads to a more random graph (Fig. 2a).
@@
[…]
@@
Metrics for performance evaluation
To quantitatively evaluate the structure of the learned latent representations, we compute a set of metrics designed to assess both local clustering and the preservation of continuous parameters.
To evaluate whether continuous graph-generating parameters vary smoothly within discrete classes, we define a class-wise Isomap gradient continuity score. For each class, we first project the latent vectors onto their first two principal components (PC1 and PC2) and embed the projected points into a one-dimensional manifold using Isomap [42] with $k=8$ nearest neighbors. This allows us to capture smooth variation even when the latent manifold is curved. We then compute the Spearman rank-order correlation coefficient [43, 44] between the resulting Isomap coordinates and the continuous parameter of interest. Taking the absolute value gives the degree of monotonic variation within the class, with a value of 1 indicating perfect monotonicity. The final score is computed as the mean absolute Spearman correlation across all classes.
For the Watts–Strogatz graphs, we compute this score with respect to the rewiring probability $p$, using the number of initial neighbors $C$ to define the classes. For graphs generated by the Vicsek model, the classes are defined by the flocking radius $R_{f}$, and the continuous parameter is the noise level $\eta$. For the brain connectivity dataset, the continuous parameter is the participants’ age, and since no discrete classes are defined, the score is computed over the entire dataset as a single group. This metric is not computed for the SMLM dataset, which lacks an associated continuous parameter.
For methods such as MIAGAE with an 8-dimensional latent space, we observe that the Isomap gradient continuity score can be low even when the AUC or $R^{2}$ values are moderate. In these cases, the structure of the latent space projection may affect the outcome of the Isomap embedding. Specifically, for the brain connectivity dataset, the first two principal components of the MIAGAE-8 latent space form a nearly circular distribution, with most of the variation captured along the first principal component. Since Isomap extracts the dominant one-dimensional manifold within each class, the resulting embedding may not align with the direction of variation associated with the continuous parameter (in this case, age), which is primarily captured by the second principal component. This misalignment leads to a reduced absolute Spearman correlation and consequently a lower continuity score.
To assess the separability of graphs in latent space based on discrete labels, we use a linear support vector machine (SVM) and evaluate classification performance using the area under the receiver operating characteristic curve (AUC). For the SMLM dataset, the task is to distinguish ring-like from spot-like structures. For the brain connectivity data, participants are divided into two age groups based on the median age (55 years). The AUC is computed using both the first two principal components and the full eight-dimensional latent representations, where applicable.
For the brain connectivity data, we further assess how much of the variance in age is captured by the latent representation by computing the coefficient of determination ($R^{2}$[40]. This is done by fitting a linear regression model with age as the target variable and either the first two principal components or all eight latent dimensions as input features.